# SpectraLift: Physics-Guided Spectral-Inversion Network for Self-Supervised Hyperspectral Image Super-Resolution

### v1

@@ Abstract
High-spatial-resolution hyperspectral images (HSI) are essential for applications such as remote sensing and medical imaging, yet HSI sensors inherently trade spatial detail for spectral richness. Fusing high-spatial-resolution multispectral images (HR-MSI) with low-spatial-resolution hyperspectral images (LR-HSI) is a promising route to recover fine spatial structures without sacrificing spectral fidelity. Most state-of-the-art methods for HSI-MSI fusion demand point spread function (PSF) calibration or ground truth high resolution HSI (HR-HSI), both of which are impractical to obtain in real world settings. We present SpectraLift, a fully self-supervised framework that fuses LR-HSI and HR-MSI inputs using only the MSI’s Spectral Response Function (SRF). SpectraLift trains a lightweight per-pixel multi-layer perceptron (MLP) network using () a synthetic low-spatial-resolution multispectral image (LR-MSI) obtained by applying the SRF to the LR-HSI as input, () the LR-HSI as the output, and () an spectral reconstruction loss between the estimated and true LR-HSI as the optimization objective. At inference, SpectraLift uses the trained network to map the HR-MSI pixel-wise into a HR-HSI estimate. SpectraLift converges in minutes, is agnostic to spatial blur and resolution, and outperforms state-of-the-art methods on PSNR, SAM, SSIM, and RMSE benchmarks.
@@ 1 Introduction
Hyperspectral imaging (HSI) captures hundreds of narrow spectral bands, enabling precise material discrimination. However, dispersive optics and photon splitting impose a trade-off: richer spectral resolution entails coarser spatial detail, which blurs small features, exacerbates mixed-pixel effects, and degrades classification accuracy. Fusion-based hyperspectral super-resolution (HSI-SR) addresses this trade-off by combining a LR-HSI with a HR-MSI of the same scene. Early approaches (pan-sharpening, Bayesian inference, matrix factorization, tensor decomposition) leverage handcrafted priors but often struggle with complex spectral structures. Recent deep learning-based methods deliver high reconstruction fidelity but depend on large labeled datasets, intricate architectures, and opaque “black-box” mappings that hinder transparency.
Unsupervised and hybrid schemes incorporate physics priors (e.g., PSF models, sparsity, low-rank constraints) to reduce supervision and improve interpretability. Yet they typically demand precise PSF knowledge – impractical given variable sensor blur – or rely on scarce ground-truth HR-HSI. State-of-the-art unsupervised methods try to estimate the PSF instead of strictly requiring precise knowledge of it. However, blurring in hyperspectral sensors is caused due to unknown and variable optical characteristics such as lens diffraction and sensor-specific aberrations, resulting in an inherent ambiguity. It is thus extremely difficult to estimate the PSF that caused the blur, which limits real-world deployment and raises concerns about model robustness and trustworthiness.
We propose SpectraLift, a self-supervised, lightweight framework that fuses LR-HSI and HR-MSI using only the multispectral sensor’s SRF. Note that the SRF is manufacturer-defined, well-documented, and routinely used for radiometric and atmospheric correction; common approximations such as Gaussian estimations also work well in practice. A synthetic LR-MSI is generated for training purposes by applying the SRF to the LR-HSI. This LR-MSI and the original LR-HSI serve (pixelwise) as the input-output pair for a compact MLP that we call the Spectral Inversion Network (SIN). The SIN learns an implicit MSI to HSI spectral mapping for the training images by minimizing an loss between the estimated and true LR-HSI. At inference, the trained SIN applies its learned spectral mapping to output a HR-HSI from a HR-MSI input.
@@
During the final preparation of this manuscript, we became aware of SSSR [SSSR], which also employs an MLP-based spectral mapping optimized with an MSE loss. While there is a similarity between our approach and SSSR in the use of an MLP for mapping MSI to HSI, crucially, unlike SSSR, SpectraLift does not rely on scene-specific PSF calibration. Instead, it leverages only the known SRF of the MSI sensor, enabling practical deployment in real-world scenarios where PSFs are unknown or difficult to estimate. We have included SSSR as one of the baselines in the experiments of Section 3. We believe that SpectraLift is the first self-supervised HSI-SR method that is fully agnostic to PSF information while achieving a more robust formulation and superior performance.
@@ 2 Proposed Method
The MSI sensor’s SRF acts as a band compression operator, mapping high-dimensional HSI spectra to lower-dimensional MSI measurements. SpectraLift, illustrated in Figure 1, formulates HSI-SR as the self-supervised, per-pixel inversion of this spectral compression. Given only a pair of LR-HSI and HR-MSI and the SRF (or its approximation), a compact MLP learns to reverse the SRF-induced degradation and reconstruct the HSI from an MSI input. SpectraLift uses only the LR-HSI image for training and the HR-MSI image for inference.
@@ 2.1 Spectral Inversion Network (SIN)
To achieve state-of-the-art performance, we tuned the scheduler parameters – specifically, the initial, maximum, and final learning rates for the One-Cycle policy, and the maximum and minimum learning rates for cosine annealing with restarts. However, SpectraLift achieves strong performance even with the default learning rate values provided in our pre-executed Jupyter notebooks. These notebooks, available in our GitHub repository https://github.com/ritikgshah/SpectraLift, contain the exact configurations used for each experiment.
@@ 2.3 SIN Insights and Limitations
Physics-Guided Interpretability: Our sole physics assumption is the SRF , which is routinely provided by MSI manufacturers. This contrasts sharply with PSF-dependent unsupervised methods (and in particular [SSSR]), since accurate PSF estimation is difficult or impossible in practice. With only a lightweight network and a clear loss, SpectraLift converges rapidly and yields directly interpretable spectral inversion filters and avoids the opaque “black-box” nature of most current state-of-the-art models.
Limitations: A limitation of SpectraLift arises when the HR-MSI input contains only a single spectral band (e.g., a monochrome image). In this extreme case, predicting a full hyperspectral signature from a scalar quantity is extremely ill-posed. In contrast, other state-of-the-art methods can exploit spatial context or impose assumptions about the similarity of spectra in neighboring pixels of the high-resolution output. While such assumptions can improve performance for highly degraded inputs, they may also reduce generalization when the input retains more spectral information. Furthermore, we note that single band HR images are rarely acquired in real world image fusion scenarios.
@@ 3 Experiments and Analysis
We evaluate SpectraLift on synthetic and real-world benchmarks. We compare against eight state-of-the-art baselines: four unsupervised (SSSR [SSSR], MIAE [MIAE], C2FF [C2FF], SDP [SDP]) and four supervised (GuidedNet [GuidedNet], FeINFN [FeINFN], FusFormer [FusFormer], MIMO-SST [MIMO]). All datasets, precomputed results, and detailed instructions for replication are available at https://github.com/ritikgshah/SpectraLift. We have made every effort to ensure reproducibility by providing pre-executed Jupyter notebooks, Python scripts for end-to-end runs, configuration files, and environment replication scripts for seamless setup. This enables reviewers and practitioners to verify results and explore extensions with minimal friction.
Quality Metrics: For synthetic data, we evaluate SpectraLift using six widely adopted metrics that capture complementary aspects of reconstruction quality in hyperspectral image super-resolution:
@@
We also profile execution time, number of parameters, inference peak GPU memory used and inference FLOPs to assess model complexity. For the inference GPU memory used and inference FLOPs we consider these for a single forward pass of the inputs through the model.
Synthetic Datasets: Following Wald’s protocol [wald], for each ground-truth HSI (GT): Washington DC Mall (DC), Kennedy Space Center (KSC), Botswana, Pavia University (Pavia U), and Pavia Center (Pavia), we generate LR-HSI by first applying spatial convolution of the GT with one of ten PSFs shown in Figure 2 (Gaussian, Kolmogorov, Airy, Moffat, Sinc, Lorentzian Squared, Hermite, Parabolic, Gabor, Delta), all with kernel size of (15,15). We then downsample the convolution output by , and finally add Gaussian white noise with SNR matched to : (4, 35 dB), (8, 30 dB), (16, 25 dB), (32, 20 dB), yielding LR-HSIs for each GT. HR-MSI are synthesized by applying SRFs for (IKONOS), (WorldView-2), and (WorldView-3), then adding 40 dB Gaussian white noise, producing 5 HR-MSIs for each GT.
@@
Tables 1-5 report mean quality and complexity metrics over these 80 LR-HSI/HR-MSI pairs for the Washington DC Mall, Kennedy Space Center, Pavia University, Pavia Center, and Botswana benchmarks, respectively, while Tables 6-10 show the same quality metrics when the single-band HR image is omitted. We do not include the model complexity metrics in tables 6-10 since they have almost no difference compared to the ones reported in tables 1-5 . Across all five synthetic datasets, SpectraLift outperforms unsupervised baselines on almost all quality measures. Supervised methods enjoyed a modest advantage – they were trained on a 75% crop of the same scene (distinct from the 25% test region), allowing them to learn scene-specific spectral behaviors. Despite lacking this advantage, on naturally occurring scenes (Washington DC Mall, Kennedy Space Center, Botswana) SpectraLift even surpasses most supervised approaches, attaining the highest PSNR (35.96dB) and lowest SAM (3.22) on DC, state-of-the-art RMSE (0.04472), PSNR (27.10dB), SSIM (0.925), ERGAS (8.41) and SAM (8.70) on KSC, and the highest SSIM (0.974) and lowest SAM (1.46) on Botswana. Only MIMO-SST edges ahead, but with a more complex method relying on ”black box” models with knowledge of the scene specific spectra. On urban scenes with more man-made materials (Pavia University, Pavia Center), scene-specific spectral information gives supervised models like MIMO-SST and FeINFN an edge, but SpectraLift still outperforms unsupervised baselines, demonstrating its advantage when comparing methods with equal training knowledge. This advantage would explain their subpar metrics on DC, KSC, and Botswana combined with their good performance against unsupervised methods on the Pavia University and Pavia Center datasets. When we omit the single-band HR image – reflecting more realistic operational scenarios – SpectraLift rises to the top on the natural scenes: it secures best RMSE, PSNR, SSIM and SAM on DC (Table 6), best results for all metrics (except UIQI) on KSC (Table 7), and best SSIM, UIQI, and SAM on Botswana (Table 10), while remaining highly competitive on the urban benchmarks. These results underscore its effectiveness when the HR input has more than a single band.
@@
Real-World Data: We further evaluate SpectraLift on the University of Houston (UH) dataset from the 2018 IEEE GRSS Data Fusion Contest [UH_data], which provides an LR-HSI () and an HR-RGB (). We extract two distinct regions from the LR-HSI and their aligned HR-RGB patches, downsample the RGB to (), and apply all unsupervised methods (including SpectraLift) to recover a HR-HSI. Testing is restricted to a small part of the UH dataset in order to ensure that baseline methods could process the inputs in a single batch; if batch processing were used, the output of each batch would have to be stitched together and visible seam artifacts would appear (as can be seen for FusFormer in Fig. 3(a) and 3(c), which could not process each region in a single batch). SpectraLift’s pixel-wise mapping is immune to such effects when inferred using batch processing.
@@
Since UH lacks HR-HSI ground truth and its PSF/SRF are unknown, we use the IKONOS RGB SRF and let each unsupervised method estimate the PSF. Supervised baselines were trained on Pavia Center HSI with inputs being generated by following Wald’s protocol for via three overlapping HR-RGB patches and the corresponding LR-HSI patches, and then evaluated on the UH test pairs. For FusFormer, 53 non-overlapping HR-RGB patches of size and the corresponding LR-HSI patches had to be used to meet memory limits.
It is important to note that we do not report proxy metrics such as Quality with No Reference (QNR) for the UH dataset. These metrics, while useful in certain controlled scenarios, are highly sensitive to the assumed PSF and SRF characteristics. In our case, no information regarding the PSF or exact SRF of the MSI acquisition is available, making such metrics unreliable and potentially misleading. Instead, we assess real-world performance through visual comparisons (Fig. 3(a), 3(c)) and spectral fidelity plots (Fig. 3(b), 3(d)).
@@
Figure 3(b) plots the spectral signatures at three manually selected pixels – car, bare earth, and highway – marked in red in Figure 3(a); Figure 3(d) does the same for concrete (residential building), metal (car), and grass pixels in Figure 3(c). Because the spatial downsampling factor is , each LR-HSI pixel corresponds to an patch in the HR grid. To compare spectra, we extract the spectrum of the LR pixel at and pair it with the spectrum of the HR-HSI at the top-right corner of its corresponding block, namely . For example, the LR pixel at maps to the HR position , so in Figures 3(b) and 3(d) we plot LR-HSI pixel (8,8) against HR-HSI pixel (64,64) for a direct spectral fidelity comparison.
SpectraLift delivers crisp, high-resolution reconstructions in both UH test scenes, whereas competing methods exhibit various spatial and color artifacts. In Figure 3(c), several baselines (MIAE, C2FF, SSSR, and FeINFN) introduce noticeable color shifts, producing HR-HSIs with unnatural tints – while SpectraLift faithfully preserves the true colors of the underlying HR-RGB. Some baselines produce HR-HSIs with spatial inconsistencies as well, such as Guided Net introducing small checkerboard artifacts that are noticeable when zoomed in. Temporal misalignment between the LR-HSI and HR-RGB is most pronounced in Figure 3(a), where moving cars occupy different positions in the two inputs. Methods that fuse the modalities jointly (MIAE, FeINFN, FusFormer, and MIMO) produce ghosting and smearing around those vehicles. Conversely, SpectraLift’s HR-HSIs are sharp, align strongly with the HR-RGB, and preserve the scene’s spatial structure.
@@
Across both UH scenes, SpectraLift’s per-pixel inversion recovers spectral shapes that mirror those of the LR-HSI, particularly for naturally occurring materials such as grass and bare earth (Figures 3(d), 3(b)). While absolute reflectance values deviate – an expected consequence of the ill-posed spectral unmixing – the overall spectral structure is faithfully preserved, supporting downstream analysis tasks that rely on relative band signatures. For certain man-made structures like highways and building rooftops, other baselines produce HR-HSIs whose spectra better aligns with that of the LR-HSI.
4 Ablation

### v2

@@ Abstract
High–spatial-resolution Hyperspectral Images (HSI) are essential for applications such as remote sensing and medical imaging, yet HSI sensors inherently trade spatial detail for spectral richness. Fusing High–spatial-Resolution Multispectral Images (HR-MSI) with Low–spatial-Resolution Hyperspectral Images (LR-HSI) is a promising route to recover fine spatial structures without sacrificing spectral fidelity. Most state of the art methods for HSI-MSI fusion demand Point Spread Function (PSF) calibration or ground truth High-spatial-Resolution HSI (HR-HSI), both of which are impractical to obtain in real world settings. We present SpectraLift, a fully self-supervised framework that fuses LR-HSI and HR-MSI inputs using only the MSI’s Spectral Response Function (SRF). SpectraLift trains a lightweight per-pixel Multi-Layer Perceptron (MLP) network using () a synthetic Low–spatial-Resolution Multispectral Image (LR-MSI) obtained by applying the SRF to the LR-HSI as input, () the LR-HSI as the target, and () an spectral reconstruction loss between the estimated and true LR-HSI as the optimization objective. At inference, SpectraLift uses the trained network to map the HR-MSI pixel-wise into a HR-HSI estimate. SpectraLift converges in minutes, is agnostic to spatial blur and resolution, and outperforms state-of-the-art methods on PSNR, SAM, SSIM, and RMSE benchmarks.
@@ 1 Introduction
Hyperspectral images (HSI) capture hundreds of narrow spectral bands, enabling precise material discrimination. However, dispersive optics and photon splitting impose a trade-off: richer spectral resolution entails coarser spatial detail, which blurs small features, exacerbates mixed-pixel effects, and degrades classification accuracy. Fusion-based Hyperspectral Super-Resolution (HSI-SR) addresses this trade-off by combining a LR-HSI with a HR-MSI of the same scene. Early approaches (pan-sharpening, Bayesian inference, matrix factorization, tensor decomposition) leverage handcrafted priors but often struggle with complex spectral structures. Recent deep learning-based methods deliver high reconstruction fidelity but depend on scarce ground-truth HR-HSI, intricate architectures, and opaque “black-box” mappings that hinder transparency.
Unsupervised and hybrid schemes incorporate physics priors (e.g., PSF models, sparsity, low-rank constraints) to reduce supervision and improve interpretability. Yet they typically demand precise PSF knowledge—impractical given variable sensor blur. State-of-the-art unsupervised methods try to estimate the PSF instead of strictly requiring precise knowledge of it. However, blurring in hyperspectral sensors is caused due to unknown and variable optical characteristics such as lens diffraction and sensor-specific aberrations, resulting in an inherent ambiguity. It is thus extremely difficult to estimate the PSF that caused the blur, which limits real-world deployment and raises concerns about model robustness and trustworthiness.
We propose SpectraLift, a self-supervised, lightweight framework that fuses LR-HSI and HR-MSI using only the multispectral sensor’s SRF. Note that the SRF is manufacturer-defined, well-documented, and routinely used for radiometric and atmospheric correction; common approximations such as Gaussian estimations also work well in practice. A synthetic LR-MSI is generated for training purposes by applying the SRF to the LR-HSI. This LR-MSI and the original LR-HSI serve (pixelwise) as the input-output pair to train a compact MLP that we call the Spectral Inversion Network (SIN). The SIN learns an implicit MSI to HSI spectral mapping for the training images by minimizing an loss between the estimated and true LR-HSI. At inference, the trained SIN applies its learned spectral mapping to output a HR-HSI from a HR-MSI input.
@@
During the preparation of this manuscript we became aware of [1], which also employs an MLP for spectral mapping that is optimized with an MSE loss. However, their training strategy generates LR-MSI by spatially blurring and downsampling HR-MSI – thus requiring precise PSF knowledge. In contrast, SpectraLift relies solely on the SRF and trains with a pixel-wise loss, yielding a more robust formulation and superior results than [1], as shown in Section 3.
@@ 2 Proposed Method
The MSI sensor’s SRF acts as a band compression operator, mapping high-dimensional HSI spectra to lower-dimensional MSI measurements. SpectraLift, illustrated in Figure 1, formulates HSI-SR as the self-supervised, per-pixel inversion of this spectral compression. Given only a pair of LR-HSI and HR-MSI, as well as the SRF (or its approximation), a compact MLP learns to reverse the SRF-induced degradation and reconstruct the HSI from an MSI input. SpectraLift uses only the LR-HSI image for training and the HR-MSI image for inference.
@@ 2.1 Spectral Inversion Network (SIN)
We tuned the scheduler parameters – specifically, the initial, maximum, and final learning rates for the One-Cycle policy, and the maximum and minimum learning rates for cosine annealing with restarts. The learning rates used in our experiments are provided in our pre-executed Jupyter notebooks. These notebooks, available in our GitHub repository (see section 3), contain the exact configurations used for each experiment.
@@ 2.3 SIN Insights and Limitations
Physics-Guided Interpretability: Our sole physics assumption is the SRF , which is routinely provided by MSI manufacturers. This contrasts sharply with PSF-dependent unsupervised methods (and in particular [1]), since accurate PSF estimation is difficult or impossible in practice. With only a lightweight network and a clear loss, SpectraLift converges rapidly and yields directly interpretable spectral inversion filters, avoiding the opaque “black-box” nature of most current state-of-the-art models.
Limitations: A limitation of SpectraLift arises when the HR-MSI input contains only a single spectral band (e.g., a monochrome image). In this extreme case, predicting a full hyperspectral signature from a scalar quantity is extremely ill-posed. In contrast, other state-of-the-art methods can exploit spatial context or impose assumptions about the similarity of spectra in neighboring pixels of the high-resolution output. While such assumptions can improve performance for highly degraded inputs, they may also reduce generalization when the input retains more spectral information. Furthermore, we note that single band HR images are rarely acquired in real world image fusion scenarios. Since SpectraLift only learns to invert the scene dependent SRF, it must be retrained for each new LR-HSI/HR-MSI pair. The SIN trained on one scene cannot be directly transferred to another, limiting the potential for cross-scene generalization which is a common limitation among state-of-the-art unsupervised/self-supervised HSI-SR methods.
@@ 3 Experiments and Analysis
We evaluate SpectraLift on synthetic and real-world benchmarks. We compare against eight state-of-the-art baselines: four unsupervised (SSSR [1], MIAE [2], C2FF [3], SDP [4]) and four supervised (GuidedNet [5], FeINFN [6], FusFormer [7], MIMO-SST [8]). All datasets, precomputed results, and detailed instructions for replication are available in our GitHub repository, which can be accessed at https://github.com/ritikgshah/SpectraLift. We have made every effort to ensure reproducibility by providing pre-executed Jupyter notebooks, Python scripts for end-to-end runs, configuration files, and environment replication scripts for seamless setup. This enables reviewers and practitioners to verify results and explore extensions with minimal friction.
3.1 Quality Metrics
For synthetic data, we evaluate SpectraLift using six widely adopted metrics that capture complementary aspects of reconstruction quality in hyperspectral image super-resolution:
@@ 3.1 Quality Metrics
We also profile execution time, number of parameters, inference peak GPU memory used and inference FLOPs to assess model complexity. For the inference GPU memory used and inference FLOPs we consider these for a single forward pass of the inputs through the model. In the tables below, we specify whether the best value of a metric is lower () or higher ().
3.2 Synthetic Datasets
Following Wald’s protocol [9], for each ground-truth HSI (GT): Washington DC Mall (DC), Kennedy Space Center (KSC), Botswana, Pavia University (Pavia U), and Pavia Center (Pavia), we generate LR-HSI by first convolving the GT with one of ten PSFs shown in Figure 2 (Gaussian, Kolmogorov, Airy, Moffat, Sinc, Lorentzian Squared, Hermite, Parabolic, Gabor, Delta), all with kernel size of (15,15). We then downsample the convolution output by , and finally add Gaussian white noise with SNR matched to : (4, 35 dB), (8, 30 dB), (16, 25 dB), (32, 20 dB), yielding LR-HSIs for each GT. HR-MSI are synthesized by applying SRFs for (IKONOS), (WorldView-2), and (WorldView-3), followed by adding Gaussian white noise with SNR=40 dB, producing 5 HR-MSIs for each GT.
@@ 3.2 Synthetic Datasets
Tables 1-5 report mean quality and complexity metrics over these 80 LR-HSI/HR-MSI pairs for the Washington DC Mall, Kennedy Space Center, Pavia University, Pavia Center, and Botswana benchmarks, respectively, while Tables 6-10 show the same quality metrics when the single-band HR image is omitted. We do not include the model complexity metrics in tables 6-10 since they have almost no difference compared to the ones reported in tables 1-5 . Across all five synthetic datasets, SpectraLift outperforms unsupervised baselines on almost all quality measures. Supervised methods enjoyed a modest advantage – they were trained on a 75% crop of the same scene (distinct from the 25% test region), allowing them to learn scene-specific spectral behaviors. Despite lacking this advantage, on naturally occurring scenes (Washington DC Mall, Kennedy Space Center, Botswana) SpectraLift even surpasses most supervised approaches, attaining the highest PSNR and lowest SAM on DC, state-of-the-art RMSE, PSNR, SSIM, ERGAS and SAM on KSC, and the highest SSIM and lowest SAM on Botswana. Only MIMO-SST edges ahead, but with a more complex method relying on ”black box” models with knowledge of the scene specific spectra. On urban scenes with more man-made materials (Pavia University, Pavia Center), scene-specific spectral information gives supervised models like MIMO-SST and FeINFN an edge, but SpectraLift still outperforms unsupervised baselines, demonstrating its advantage when comparing methods with equal training knowledge. When we omit the single-band HR image – reflecting more realistic operational scenarios – SpectraLift rises to the top on the natural scenes: it secures best RMSE, PSNR, SSIM and SAM on DC (Table 6), best results for all metrics (except UIQI) on KSC (Table 7), and best SSIM, UIQI, and SAM on Botswana (Table 10), while remaining highly competitive on the urban benchmarks. These results underscore its effectiveness when the HR input has more than a single band.
@@ 3.3 Real-World Data
We further evaluate SpectraLift on the University of Houston (UH) dataset from the 2018 IEEE GRSS Data Fusion Contest [10], which provides an LR-HSI () and an HR-RGB (). We remove the last 2 bands from the LR-HSI due to noise, resulting in an LR-HSI of shape (). We extract two distinct regions from the LR-HSI and their aligned HR-RGB patches, downsample the RGB to (), and apply all unsupervised methods (including SpectraLift) to recover a HR-HSI. Testing is restricted to a small part of the UH dataset and the HR-RGB is downsampled from to , in order to ensure that baseline methods could process the inputs in a single batch; if batch processing were used, the output of each batch would have to be stitched together and visible seam artifacts would appear (as can be seen for FusFormer in Fig. 3(a) and 3(c), which could not process each region in a single batch).
@@
Since UH lacks HR-HSI ground truth and its PSF/SRF are unknown, we use the IKONOS RGB SRF and let each unsupervised method estimate the PSF. Supervised baselines were trained on Pavia Center HSI with inputs being generated by following Wald’s protocol for via three overlapping HR-RGB patches and the corresponding LR-HSI patches, and then evaluated on the UH test pairs. For FusFormer, 53 non-overlapping HR-RGB patches of size and the corresponding LR-HSI patches had to be used to meet memory limits. Since no ground truth is available for this dataset and the true SRF, PSF is unknown, we avoid relying on QNR scores which are highly sensitive to these assumptions. We instead assess real-world performance via visual comparisons, generated by picking (Red = 659nm, Green = 561nm, Blue = 492nm) bands from the HR-HSI outputs of each method (Figs. 3(a) and 3(c)), and spectral fidelity plots (Figs. 3(b) and 3(d)).
@@
Figure 3(b) plots the spectral signatures at three manually selected pixels – highway, bare earth, and car pixels – marked in red in Figure 3(a); Figure 3(d) does the same for grass, metal (car), and concrete (residential building) in Figure 3(c). With a spatial downsampling factor of , each LR-HSI pixel corresponds to an region in the HR grid. To compare spectra, we pair the spectrum of the LR-HSI pixel at with that of the HR-HSI pixel at the top-right corner of the corresponding block, . For instance, in Figures 3(b) and 3(d), we compare the spectra of LR-HSI pixel and HR-HSI pixel to assess spectral fidelity.
Because the proprietary SRF of the RGB camera used to capture the HR-RGB images is unavailable, we adopt the IKONOS RGB SRF in our experiments. However, the IKONOS SRF is likely a poor approximation of the true sensor response, which remains unknown and cannot be reliably estimated without manufacturer-provided specifications. In practice, the true SRF is always known to the camera manufacturer and would be available for real-world missions, eliminating this limitation. Unfortunately, the manufacturer has not disclosed the SRF of this particular camera, preventing a more accurate approximation. This substitution introduces color shifts in the HR-HSI outputs, particularly affecting unsupervised methods that explicitly rely on the SRF. Supervised methods are also influenced, as their synthetic training data uses the same IKONOS SRF to generate LR-HSI pairs. If the true SRF were available—or if its characteristics (e.g., center, minimum, and maximum wavelengths of each band) were known, enabling a Gaussian or parametric approximation—this limitation could be alleviated.
These tints are most visible in GuidedNet, FeINFN, FusFormer, and MIMO-SST in Figure 3(a), and in MIAE, SDP, SSSR, and GuidedNet in Figure 3(c). While SpectraLift is likewise affected— evidenced by red regions appearing grey (e.g., roofs, car, truck, and traffic cones)—it better preserves spatial structure and overall chromaticity than other baselines. Additional artifacts include checkerboard patterns in GuidedNet and patch-seam artifacts in FusFormer. Temporal misalignment between LR-HSI and HR-RGB, evident in Figure 3(a) (e.g., moving cars), causes ghosting in joint-fusion methods (MIAE, FeINFN, FusFormer, MIMO-SST), whereas SpectraLift remains sharp and spatially coherent.
@@
Across both UH scenes, SpectraLift’s per-pixel inversion recovers spectral shapes closely matching those of the LR-HSI, particularly for natural materials such as grass and bare earth (Figs. 3(b), 3(d)). Although absolute reflectance values deviate—an expected outcome of the ill-posed spectral unmixing—the overall spectral structure remains well preserved, supporting downstream analysis. For certain man-made surfaces, such as highways and rooftops, other baselines achieve closer alignment with the LR-HSI spectra.
3.4 Ablation